benji raises 39 dollars by selling 5 rolls of red wrapping paper and 2 rolls of foil wrapping paper, joel raises 57 dollars by selling 3 rolls of red wrapping paper and 6 rolls of foil wrapping paper. for how much are benji and joel selling each toll of red and foil wrapping paper

Answer by JulietG(1812)

5R + 2F = 39
3R + 6F = 57
Add
8R + 8F = 96
.
There is more than one possibility here. Let's assume whole numbers only.
Red could be a dollar. That makes foil 96-8/8 or 11 each.
If R = 2, then F = 10 (8*2) + (8*10) = 96
If R = 3, then F = 9 (8*3) + (8*4) = 96
If R = 4, then F = 8
R = 5, then F = 7
R = 6, then F = 6
R = 7, then F = 5
R = 8, then F = 4
R = 9, then F = 3
R = 10, then F = 2
R = 11, then F = 1